find the four numbers in G.P. in which third term is greater than first by 9, second term is greater than foutht by 18 ? 

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Raju

let the numbers are a, ar, ar2, ar3

from the given conditions

ar2-a=9   ........(1)

ar-ar3=18   ......(2)

dividing (2) by (1)

(r-r3)/(r2-1)=18/9

r(1-r2)/(r2-1)=2

r= -2

putting r=-2 in (1)

a=3

hence numbers will be 3,-6,12,-24

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

vikas askiitian expert
13 years ago

let the numbers are   a/r  ,  a  ,ar  , ar

 

third term is greater than first by 9 so


ar=a/r +9 or  ar2 =a+9r...........1


also second term is greater than fourth by 18 so


a =ar2 +18.......2


solving 1 and 2 we get r=-2 & a=-6


therefore numbers are  3 , -6 ,12 ,-24